OpenStudy (lena772):

Find the distance traveled in 15 seconds by an object moving with a velocity of v(t)=20+7cos t feet per second?

OpenStudy (lena772):

Is 301.81ft the right answer?

OpenStudy (lena772):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

can you show me your steps .@Lena772

OpenStudy (cggurumanjunath):

@Lena772

OpenStudy (lena772):

yes i have (20+7cos(t))dt

OpenStudy (cggurumanjunath):

ok

OpenStudy (cggurumanjunath):

then

OpenStudy (lena772):

then v(t)=ds/dt

OpenStudy (cggurumanjunath):

yes correct

OpenStudy (lena772):

then 20+7cost=[20t+7sint] (t=15)

OpenStudy (lena772):

s(15)=20(15)+7sin15

OpenStudy (cggurumanjunath):

very good

OpenStudy (lena772):

s(15)=300+7(0.2589)

OpenStudy (lena772):

s(15)=300+1.81

OpenStudy (lena772):

s(15)=301.81ft?

OpenStudy (lena772):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

there was power cut so !

OpenStudy (cggurumanjunath):

@Lena772

OpenStudy (cggurumanjunath):

u are correct.

OpenStudy (lena772):

Thank you and I'm glad it's back on c:

OpenStudy (cggurumanjunath):

welcome.